# Video-based Automatic Lameness Detection of Dairy Cows using Pose Estimation and Multiple Locomotion Traits

### v1

@@ 3.1.2 Keypoint detection
We trained T-LEAP to predict the location of 9 keypoints. They represented the location of the following anatomical landmarks: Nose, Forehead, Withers, Sacrum, Caudal thoracic vertebrae, and the four Hooves (Figure 2). The location of these nine keypoints was needed for extracting the gait features described in subsection 3.2. T-LEAP was trained with sequences of 2 consecutive frames as input because the authors reported the best performance with T=2 russello_t-leap_2021 .
A pose estimation dataset was created for training and evaluating T-LEAP, using 28 videos of unique cows randomly selected from of the 272 available videos. The coordinates of the nine keypoints were annotated for each frame of the 28 videos and divided into 968 non-overlapping sequences of 2 frames. We refer to each set of consecutive frames as a sample. T-LEAP was trained with a random subset of 80% of the samples (i.e., 774 training samples) and evaluated on the remaining 20% of the samples (i.e., 194 test samples). We used the same training procedure and hyper-parameters settings as described in the original T-LEAP paper russello_t-leap_2021 .
@@
Figure 2: The 9 keypoints (anatomical landmarks) as described in russello_t-leap_2021 . The keypoints are named as follows: 1: Left-hind hoof, 2: Right-hind hoof, 3: Left-front hoof, 4: Right-front hoof 5: Nose, 6: Forehead, 7: Withers, 8: Sacrum, 9: Caudal thoracic vertebrae.
@@ 3.1.3 Keypoint correction
In our set of 272 videos, we identified 98 individual cows. There were 28 videos of unique cows included in training the pose estimation model, and thus 70 cows that the pose estimation model did not see. In their generalization experiment, the authors of T-LEAP reported a percentage of correct keypoints (PCKh@0.2) of $93.8\%$ on known cows (i.e., cows included in the training set) and a performance of $87.6\%$ on unknown cows (i.e., cows not included in the training set). It was, therefore, expected to have errors in the predicted keypoint trajectories. To deal with that, we developed a method for correcting the keypoints. First, to identify and correct large outliers in the trajectories, we used a Median-Absolute-Deviation (MAD) filter with a temporal window of size 3. We then applied a Savitzky–Golay filter savitzky1964smoothing (window=10, order=3) to smooth the trajectories temporally. Figure 3 shows examples of trajectories with outliers before and after applying the filters.
@@ 3.2 Gait features extraction
Using the keypoint trajectories, we computed six locomotion traits that were shown to be correlated with locomotion scores schlageter-tello_relation_2015 , namely Back Posture Measurement (BPM), Head Bobbing Amplitude (HBA), Tracking distance (TRK), Stride Length (STL), Stance Duration (STD) and Swing Duration (SWD). All features relied on step detection, that is, knowing when each hoof was moving (swing phase) or remained still (stance phase). Hence, in the following paragraphs, we first describe the implementation of the step detection, followed by the implementation of the gait features.
@@ 3.2.2 Step correction
The step detection was automatically controlled and corrected using the following procedure: for any given leg, mid-swings must happen before or after the stance phases, and the mid-swings must happen during the supporting phase of the opposite leg (left-right). When the step detection failed to meet these requirements, this indicated that the keypoint predictions were too noisy on that hoof. The frames with problematic steps were then removed from the keypoint trajectories, resulting in trajectories with one or several gaps. The trajectories were then trimmed to the part with the most remaining frames. Using this method, only four videos were found to have problematic step detection, and only one of them had less than two stance phases per leg. The latter was then discarded from the dataset, as at least two stance phases are needed to compute some of the features. As a result, the final dataset included 271 videos.
@@ 3.2.3 Back posture measurement (BPM)
To estimate the back posture, or curvature of the back, a similar approach as described in poursaberi_real-time_2010 was taken. A circle was fitted through the three keypoints on the spine. The curvature of a circle can be found by taking the inverse of its radius. The radius ($r$) of the fitted circle was normalized with the head length ($h$) of the cow (in pixels), as the length of cows can differ. The head length was taken as the Euclidean distance between the keypoints on the forehead and the nose. The BPM was then calculated as follows:
@@ 3.2.4 Head bobbing amplitude (HBA)
Head bobbing is defined as an exaggerated movement of the head when an affected limb lands and lifts from the ground schlageter-tello_relation_2015 ; blackie_associations_2013 . Hence, in the presence of head bobbing, the head moves significantly up and down cyclically (at least once per gait cycle). Sound subjects are expected to have a more steady head stance. Examples of a noticeable head bob and steady head stance are shown in Figure 5. The amplitude of the vertical movement (y-signal) of the forehead keypoint was used as a measure of head bobbing. The amplitude of the y-signal was computed with fast Fourier transforms cooley1965algorithm as follows: let $N_{v}$ be the number of frames in a video, let $N_{g}$ be the number of frames per gait cycle in a video, $k\in[1,N_{v}]$ the frequency, $X$ the Fourier transform of the signal, and $A_{k}$ the amplitude at frequency $k$. The value of the HBA was then assigned as the largest amplitude in a gait cycle:
@@ 3.2.5 Tracking distance (TRK)
[…]
@@ 3.2.6 Stride length difference (STL)
[…]
@@ 3.2.7 Stance duration difference (STD)
[…]
$t_{s}$ $=b_{s}-a_{s}$ (7)
@@ 3.2.8 Swing duration difference (SWD)
[…]
$w_{s}$ $=b_{s}-a_{s}$ (9)
@@
[…]
Figure 6: Distribution of the features per lameness score, where 0 corresponds to healthy, and 1 to lame.
@@
3.3 Gait classification
The layout of our machine-learning experiments is described in the next paragraphs. We first split the data into training and validation sets using cross-validation. We then trained and evaluated different classifiers to score the gait using all the extracted features. Lastly, we investigated the importance of features on classification performance.
@@ 3.3.1 Data preparation
Considering the relatively small dataset size (271 videos), the dataset was split into training and validation sets using a 5-fold cross-validation (CV) with stratified grouping. In order to prevent data leakage, the grouping was performed on the cow IDs to ensure that, in each fold, there was no overlap of cow IDs between the training and the validation set. Given this non-overlapping constraint, the stratification creates folds that retain, as much as possible, the same class distribution buitinck2013api . To ensure a balanced class distribution during training, we applied the Synthetic Minority Oversampling Technique (SMOTE) chawla2002smote to the minority classes in the training sets. SMOTE generates new training samples whose feature values are close to the other samples in the minority class. Lastly, the features were re-scaled as machine-learning models often require the features to be on a similar scale. The range of the features was re-scaled using Robust Scaling scikit-learn , which uses statistics that are robust to outliers for scaling the data.
@@ 3.3.2 Classification models
We compared the performance of the following six classifiers: Logistic Regression (LR), Random Forest (RF), Support Vector with a linear kernel (SVL) and with a radial kernel (SVR), Multi-Layer Perceptron (MLP) and Gradient Boosting Machines (GB). These classifiers were selected as they showed good performance in previous research on lameness detection zhao_automatic_2023 ; zheng_cows_2023 ; arazo_segmentation_2022 . We used a flat cross-validation approach to tune the hyper-parameters and train the models, as it is computationally less expensive than nested cross-validation, and generally results in the selection of an algorithm of similar quality to that selected via nested cross-validation wainer2021nested . The hyper-parameters of the classifiers were first optimized using a random cross-validated search of 100 iterations over the 5-folds. The classifiers were then re-trained on the 5-folds with the best set of hyper-parameters.
@@ 3.3.4 Feature importance
An additional experiment was run to investigate whether including multiple features could lead to improvements in gait scoring. The predictive value of a feature was evaluated by measuring the feature importance, that is, how much a feature contributed to a correct classification. To measure the feature importance, we selected the permutation importance method breiman2001random as it can be applied to any classifier. The importance of features was evaluated on the best-performing classifier among the 6 classifiers that were trained with all the features. The permutation importance method was performed as follows: For each cross-validation fold, the model was fitted on the training dataset and evaluated on the F1-score on the validation set. Then, a feature column from the validation set was randomly shuffled, and the model was evaluated again. The importance score was then the difference between the F1-score on the non-shuffled and the shuffled validation data. The permutations were repeated 100 times for each feature. The features were then ranked in the order of their mean importance score. To estimate whether including multiple features could lead to improvements in the gait scoring, the classifier was then retrained with the most important feature, the two most important features, and so on, gradually adding one feature in the order of their importance.
@@ 4.1 Pose estimation
The test results of T-LEAP are presented in Table 6. On average, there were 99.6% of correctly detected keypoints (PCKh@0.2). In other words, the Euclidian distance between the predicted keypoint and its ground truth was smaller than 20% of the head length in 99.6% of the cases. This is in line with the results presented in the original paper russello_t-leap_2021 , where they achieved a 99.0% detection rate on the same model with 17 keypoints. The keypoint correction and filtering were run on all 272 videos, and the MAD filter (of window size 3) identified 0.21% of outlier keypoints, whose coordinates were then corrected to the median value of the temporal window. Because of the lack of keypoint annotations on all videos, the keypoint correction could only be assessed qualitatively. The trajectories of the keypoints before and after the filtering were plotted for each video and controlled visually. The quality of the filtered trajectories was deemed balanced, in that most of the outliers could be corrected and the trajectories appeared smooth, without over-correction or flattening. The outliers that could not be corrected sufficiently led to a wrong step detection. These steps were then discarded from trajectories, as detailed in section 3.2.
@@
4.2 Gait score classification
The classification results of the different classifiers are listed in Table 7. The SVM with radial kernel, Random Forests, and Gradient Boosting classifiers performed best, with an accuracy above 79%. SVM-R had a higher specificity, while the Random Forests and Gradient Boosting had a higher sensitivity. The logistic regression, the SVM with linear kernel, and the Multi-Layer Perceptron performed slightly worse.
Table 7: Results of the classifiers using all the features. Values are expressed in %. The best results are highlighted in bold.
@@ 5.1 Video processing
The video processing consisted of the following steps: using Faster-R-CNN to detect and isolate the cows from the video frames, using T-LEAP to extract time-series of keypoint locations, and using the MAD and Savitzky–Golay filters to reduce noise from the keypoint predictions. For our set of videos, the pre-trained Faster-R-CNN worked out of the box and detected the location of the cows in each video frame. The performance of T-LEAP was on par with the results described in the original paper russello_t-leap_2021 , and it would require little effort to be transferred to videos recorded in new farms, as taghavi2023cow showed that little new training data was needed to fine-tune the T-LEAP model. However, some keypoint mis-detections needed to be corrected. The parameters for the MAD outlier filter and the smoothing Savitzky–Golay filter had to be tuned manually until a good trade-off was found between under- and over-correction. With no or insufficient correction of the keypoint trajectories, the features could give erroneous values. While with over-correction, one would run the risk of removing the true signal of keypoint trajectories, and the extracted features wouldn’t be discriminatory. For instance, if the signal of the forehead would be too flattened, the head bobbing would be systematically missed.
The videos were selected such that there was only one cow at a time in the field of view. This constraint makes the gait analysis more reliable in two ways. First, having a single cow in the field of view ensures that the cows don’t occlude each other’s body parts, making the pose estimation more reliable. Second, a single cow in the field of view ensures enough space between the cows such that they can walk at their own pace and display a voluntary gait. In practice, this constraint could be implemented by skipping the videos where the Faster-R-CNN (or any other object detector) detects more than one cow, or as done in barney2023deep , by implementing a tracking algorithm that follows each cow through the video.
@@ 5.2 Locomotion scoring
A classifier learns to classify samples from a set of labeled examples, also known as ground-truth or golden-standard. Because a classifier can only be as accurate as its golden-standard schlageter-tello_relation_2015 , a reliable locomotion scale is necessary. Here, the initial inter- and intra-observer reliability was under par. It is worth noting that the reliability is usually lower in homogeneous data because the probability of agreement by chance is higher when scores are not equally distributed schlageter-tello_effect_2014 . It is unlikely that scoring from live observations instead of from videos would have improved the scores, as schlageter2015comparison showed no difference in the reliability of inexperienced observers between live and video scoring and showed improved reliability of experienced observers when scoring from video. The quality of the ground-truth could perhaps have been further improved by organizing additional locomotion scoring sessions or by having shorter scoring sessions over multiple days. However, given that the availability of the observers was limited and that a perfect golden standard was not necessary nor likely achievable, we took other steps to address the problem of low reliability and agreement. First, because we had multiple observers, we could discard the votes from the least reliable observers. Second, we addressed the problem of class (score) imbalance by merging the levels of the scale to a binary score: normal and lame. By doing so, we then increased the quality of our golden standard to an acceptable level for running the experiments.
@@
5.3 Gait score classification
The gait-score classification task was binary (normal vs. lame) and therefore focused on lameness detection rather than fine-grained gait scoring. Fine-grained locomotion scoring is left for future research as it would require collecting more video footage with sufficient examples of gait scores of 3 and above.
The performance of the linear classifiers (i.e., logistic regression and SVM with linear kernel) was lower than the performance of the non-linear classifiers. This implies that when combining all the features, the decision boundary between the normal and lame classes is non-linear. The Multi-Layer Perceptron didn’t perform as well as the other non-linear classifiers, most likely because of the relatively small dataset. The performance of the three best classifiers SVM-R, RF, and BG, aligns with the conclusions of wainer2016comparison and wainer2021nested : they found these three binary classifiers to perform the best on 115 open-source datasets tackling a variety of real-world problems in medicine and biology (but not related to lameness detection). Although, on this dataset, the SVM classifier with radial kernel achieved the best performance in terms of accuracy and F1-score, it might not be the case for other datasets. This is a well-known machine-learning challenge, also known as the “no-free-lunch” theorem, that suggests that no algorithm can outperform all others for all problems wolpert1996lack . Our recommendation would then be to try several classifiers, and the SVMs with radial kernel, random forests, and gradient boosting classifiers provide a good starting point.

### v2

@@ 3.1.2 Keypoint detection
We trained T-LEAP to predict the location of 9 keypoints. They represented the location of the following anatomical landmarks: Nose, Forehead, Withers, Sacrum, Caudal thoracic vertebrae, and the four Hooves (Figure 2). The location of these nine keypoints was needed for extracting the gait features described in subsection 3.2. T-LEAP was trained with sequences of 2 consecutive frames as input because the authors reported the best performance with T=2 [16].
A pose estimation dataset was created for training and evaluating T-LEAP, using 28 videos of unique cows randomly selected from of the 272 available videos. The coordinates of the nine keypoints were annotated for each frame of the 28 videos and divided into 968 non-overlapping sequences of 2 frames. We refer to each set of consecutive frames as a sample. T-LEAP was trained with a random subset of 80% of the samples (i.e., 774 training samples) and evaluated on the remaining 20% of the samples (i.e., 194 test samples). We used the same training procedure and hyper-parameters settings as described in the original T-LEAP paper [16].
@@
Figure 2: The 9 keypoints (anatomical landmarks) as described in [16]. The keypoints are named as follows: 1: Left-hind hoof, 2: Right-hind hoof, 3: Left-front hoof, 4: Right-front hoof 5: Nose, 6: Forehead, 7: Withers, 8: Caudal thoracic vertebrae, 9: Sacrum.
@@ 3.1.3 Keypoint correction
In our set of 272 videos, we identified 98 individual cows. There were 28 videos of unique cows included in training the pose estimation model, and thus 70 cows that the pose estimation model did not see. In their generalization experiment, the authors of T-LEAP reported a percentage of correct keypoints (PCKh@0.2) of $93.8\%$ on known cows (i.e., cows included in the training set) and a performance of $87.6\%$ on unknown cows (i.e., cows not included in the training set). It was, therefore, expected to have errors in the predicted keypoint trajectories. To deal with that, we developed a method for correcting the keypoints. First, to identify and correct large outliers in the trajectories, we used a Median-Absolute-Deviation (MAD) filter with a temporal window of size 3. We then applied a Savitzky–Golay filter [32] (window=10, order=3) to smooth the trajectories temporally. Figure 3 shows examples of trajectories with outliers before and after applying the filters.
@@ 3.2 Gait features extraction
Using the keypoint trajectories, we computed six locomotion traits that were shown to be correlated with locomotion scores [25], namely Back Posture Measurement (BPM), Head Bobbing Amplitude (HBA), Tracking distance (TRK), Stride Length (STL), Stance Duration (STD) and Swing Duration (SWD). All features relied on step detection, that is, knowing when each hoof was moving (swing phase) or remained still (stance phase). Hence, in the following paragraphs, we first describe the implementation of the step detection, followed by the implementation of the gait features.
@@ 3.2.2 Step correction
The step detection was automatically controlled and corrected using the following procedure: for any given leg, mid-swings must happen before or after the stance phases, and the mid-swings must happen during the supporting phase of the opposite leg (left-right). When the step detection failed to meet these requirements, this indicated that the keypoint predictions were too noisy on that hoof. Only four videos were found to have problematic step detection. The frames with problematic steps were then removed from the keypoint trajectories, resulting in trajectories with one or several gaps. The trajectories were then trimmed to the part with the most remaining frames.
@@ 3.2.3 Back posture measurement (BPM)
To estimate the back posture, or curvature of the back, a similar approach as described in [6] was taken. A circle was fitted through the three keypoints on the spine. The curvature of a circle can be found by taking the inverse of its radius. The radius ($r$) of the fitted circle was normalized with the head length ($h$) of the cow (in pixels), as the length of cows can differ. The head length was taken as the Euclidean distance between the keypoints on the forehead and the nose. The BPM was then calculated as follows:
@@ 3.2.4 Head bobbing amplitude (HBA)
Head bobbing is defined as an exaggerated movement of the head when an affected limb lands and lifts from the ground [25, 9]. Hence, in the presence of head bobbing, the head moves significantly up and down cyclically (at least once per gait cycle). Sound subjects are expected to have a more steady head stance. Examples of a noticeable head bob and steady head stance are shown in Figure 5. The amplitude of the vertical movement (y-signal) of the forehead keypoint was used as a measure of head bobbing. The amplitude of the y-signal was computed with fast Fourier transforms [33] as follows: let $N_{v}$ be the number of frames in a video, let $N_{g}$ be the number of frames per gait cycle in a video, $k\in[1,N_{v}]$ the frequency, $X$ the Fourier transform of the signal, and $A_{k}$ the amplitude at frequency $k$. The value of the HBA was then assigned as the largest amplitude in a gait cycle:
@@ 3.2.5 Tracking distance (TRK)
The tracking distance is defined as the horizontal distance (x-coordinate) between the landing position of the front hoof and the subsequent landing position of the hind hoof of the same side. If the hind hoof lands at the same location as the front hoof, it indicates no serious walking problem [5], and the TRK value is equal (or close) to 0. The tracking distance was measured on the left (TRKL) and right (TRKR) side of the cow and was normalized to the head length ($h$) as follows: for any given side (left, right), let $x_{f}$ and $x_{h}$ be the x-coordinates of the front and hind hooves, Let $s$ be the start frame of a stance phase on the front hoof, and $s+1$ the start frame of the subsequent stance phase on the hind hoof. When there was more than one value per side, the median TRK value of that side was returned.
@@ 3.2.6 Stride length difference (STL)
The stride length is defined as the horizontal distance between two successive landings of the same hoof. The stride length ($l$) was measured for each hoof between each successive stance phase ($s$) and normalized to the head length ($h$). If there was more than one stride length per hoof, the median value was kept. We measured the difference in stride length between the left and right sides for the hind (STLH) and front (STLF) legs as follows:
@@ 3.2.7 Stance duration difference (STD)
We define the stance duration as the time (in seconds) between the start ($a$) and end ($b$) of each stance phase. The time in seconds was derived from the frame rate of the video recording (here, 30 fps).
The stance duration ($t$) was measured per hoof for each stance phase. If a leg had more than one stance phase, the median duration was used. We measured the difference in duration between the left and right sides for the hind (STDH) and front (STDF) legs as follows:
$t$ $=b-a$ (7)
@@ 3.2.8 Swing duration difference (SWD)
We define the swing duration as the time (in seconds) between the ($a$) and end ($b$) of each swing phase. The time in seconds was derived from the frame rate of the video recording (here, 30 fps).
The swing duration ($w$) was measured per hoof for each swing phase. If a leg had more than one swing phase, the median duration was used. We measured the difference in duration between the left and right sides for the hind (SWDH) and front (SWDF) legs as follows:
$w$ $=b-a$ (9)
@@
TRKL Tracking distance on the left side
TRKR Tracking distance on the right side
STLF Stride length difference between left- and right-front hooves
STLH Stride length difference between left- and right-hind hooves
STDF Stance duration difference between left- and right-front hooves
STDH Stance duration difference between left- and right-hind hooves
SWDF Swing duration difference between left- and right-front hooves
SWDH Swing duration difference between left- and right-hind hooves
Figure 6: Distribution of the features per lameness class, where 0 corresponds to normal, and 1 to lame.
@@
3.3 Lameness classification
The layout of our machine-learning experiments is described in the next paragraphs. We first split the data into training and validation sets using cross-validation. We then trained and evaluated different binary classifiers to detect lameness using all the extracted features. Lastly, we investigated the importance of features on classification performance.
@@ 3.3.1 Data preparation
Considering the relatively small dataset size (272 videos), the dataset was split into training and validation sets using a 5-fold cross-validation (CV) with stratified grouping. In order to prevent data leakage, the grouping was performed on the cow IDs to ensure that, in each fold, there was no overlap of cow IDs between the training and the validation set. Given this non-overlapping constraint, the stratification creates folds that retain, as much as possible, the same class distribution [34]. To ensure a balanced class distribution during training, we applied the Synthetic Minority Oversampling Technique (SMOTE) [35] to the minority classes in the training sets. SMOTE generates new training samples whose feature values are close to the other samples in the minority class. Lastly, the features were re-scaled as machine-learning models often require the features to be on a similar scale. The range of the features was re-scaled using Robust Scaling [36], which uses statistics that are robust to outliers for scaling the data.
@@ 3.3.2 Classification models
We compared the performance of the following six classifiers: Logistic Regression (LR), Random Forest (RF), Support Vector with a linear kernel (SVL) and with a radial kernel (SVR), Multi-Layer Perceptron (MLP) and Gradient Boosting Machines (GB). These classifiers were selected as they showed good performance in previous research on lameness detection [8, 7, 14]. We used a flat cross-validation approach to tune the hyper-parameters and train the models, as it is computationally less expensive than nested cross-validation, and generally results in the selection of an algorithm of similar quality to that selected via nested cross-validation [37]. The hyper-parameters of the classifiers were first optimized using a random cross-validated search of 100 iterations over the 5-folds. The classifiers were then re-trained on the 5-folds with the best set of hyper-parameters.
@@ 3.3.4 Feature importance
An additional experiment was run to investigate whether including multiple features could lead to improvements in lameness classification. The predictive value of a feature was evaluated by measuring the feature importance, that is, how much a feature contributed to a correct classification. To measure the feature importance, we selected the permutation importance method [38] as it can be applied to any classifier. The importance of features was evaluated on the best-performing classifier among the 6 classifiers that were trained with all the features. The permutation importance method was performed as follows: For each cross-validation fold, the model was fitted on the training dataset and evaluated on the F1-score on the validation set. Then, a feature column from the validation set was randomly shuffled, and the model was evaluated again. The importance score was then the difference between the F1-score on the non-shuffled and the shuffled validation data. The permutations were repeated 100 times for each feature. The features were then ranked in the order of their mean importance score. To estimate whether including multiple features could lead to improvements in the lameness classification, the classifier was then retrained with the most important feature, the two most important features, and so on, gradually adding one feature in the order of their importance.
@@ 4.1 Pose estimation
The test results of T-LEAP are presented in Table 4. On average, there were 99.6% of correctly detected keypoints (PCKh@0.2). In other words, the Euclidean distance between the predicted keypoint and its ground truth was smaller than 20% of the head length in 99.6% of the cases. This is in line with the results presented in the original paper [16], where they achieved a 99.0% detection rate on the same model with 17 keypoints. The keypoint correction and filtering were run on all 272 videos, and the MAD filter (of window size 3) identified 0.21% of outlier keypoints, whose coordinates were then corrected to the median value of the temporal window. Because of the lack of keypoint annotations on all videos, the keypoint correction could only be assessed qualitatively. The trajectories of the keypoints before and after the filtering were plotted for each video and controlled visually. The quality of the filtered trajectories was deemed balanced, in that most of the outliers could be corrected and the trajectories appeared smooth, without over-correction or flattening. The outliers that could not be corrected sufficiently led to a wrong step detection. These steps were then discarded from trajectories, as detailed in section 3.2.
@@
4.2 Lameness detection
The results of the different binary classifiers are listed in Table 5. The SVM with radial kernel, Random Forests, and Gradient Boosting classifiers performed best, with an accuracy above 79%. SVM-R had a higher specificity, while the Random Forests and Gradient Boosting had a higher sensitivity. The logistic regression, the SVM with linear kernel, and the Multi-Layer Perceptron performed slightly worse.
Table 5: Results of the binary classifiers using all the features. Values are expressed in %. The best results are highlighted in bold.
@@ 5.1 Video processing
The video processing consisted of the following steps: using Faster-R-CNN to detect and isolate the cows from the video frames, using T-LEAP to extract time-series of keypoint locations, and using the MAD and Savitzky–Golay filters to reduce noise from the keypoint predictions. For our set of videos, the pre-trained Faster-R-CNN worked out of the box and detected the location of the cows in each video frame. The performance of T-LEAP was on par with the results described in the original paper [16], and it would require little effort to be transferred to videos recorded in new farms, as [18] showed that little new training data was needed to fine-tune the T-LEAP model. However, some keypoint mis-detections needed to be corrected. The parameters for the MAD outlier filter and the smoothing Savitzky–Golay filter had to be tuned manually until a good trade-off between under- and overcorrection was found. With no or insufficient correction of the keypoint trajectories, the features could give erroneous values. While with over-correction, one would run the risk of removing the true signal of keypoint trajectories, and the extracted features wouldn’t be discriminatory. For instance, if the signal of the forehead would be too flattened, the head bobbing would be systematically missed.
The videos were selected such that there was only one cow at a time in the field of view. This constraint makes the gait analysis more reliable in two ways. First, having a single cow in the field of view ensures that the cows don’t occlude each other’s body parts, making the pose estimation more reliable. Second, a single cow in the field of view ensures enough space between the cows such that they can walk at their own pace and display a voluntary gait. In practice, this constraint could be implemented by skipping the videos where the Faster-R-CNN (or any other object detector) detects more than one cow, or as done in [17], by implementing a tracking algorithm that follows each cow through the video.
Another constraint for selecting the video was that the cows had to walk from left to right. However, our proposed method would still work if the cows were walking in the opposite direction. We can determine the walking direction by examining the horizontal movement of the keypoints along the x-axis. The values of the keypoints x-coordinates increase over time when the cows walk to the right and decrease when they walk to the left. Therefore, if the cows were to walk from right to left, their keypoint trajectories could simply be mirrored on the x-axis before extracting the features.
@@ 5.2 Locomotion scoring
A classifier learns to classify samples from a set of labeled examples, also known as ground-truth or golden-standard. Because a classifier can only be as accurate as its golden-standard [25], a reliable locomotion scale is necessary. Here, the initial inter- and intra-observer reliability was under par. It is worth noting that the reliability is usually lower in homogeneous data because the probability of agreement by chance is higher when scores are not equally distributed [24]. It is unlikely that scoring from live observations instead of from videos would have improved the scores, as [39] showed no difference in the reliability of inexperienced observers between live and video scoring and showed improved reliability of experienced observers when scoring from video. The quality of the ground-truth could perhaps have been further improved by organizing additional locomotion scoring sessions or by having shorter scoring sessions over multiple days. However, given that the availability of the observers was limited and that a perfect golden standard was not necessary nor likely achievable, we took other steps to address the problem of low reliability and agreement. Firstly, because we had multiple observers, we could discard the votes from the least reliable observers. Secondly, we addressed the problem of class (score) imbalance by merging the levels of the scale to a binary score: normal and lame. Although these steps improved the quality of our golden standard, some biases might have remained, which could limit the classifiers’ accuracy.
@@
5.3 Lameness detection
The lameness detection was performed as a binary classification task (normal vs. lame) and therefore focused on lameness detection rather than fine-grained gait scoring. Fine-grained locomotion scoring is left for future research as it would require collecting more video footage with sufficient examples of gait scores of 3 and above.
The performance of the linear classifiers (i.e., logistic regression and SVM with linear kernel) was lower than that of the non-linear classifiers. This implies that when combining all the features, the decision boundary between the normal and lame classes is non-linear. The Multi-Layer Perceptron didn’t perform as well as the other non-linear classifiers, most likely because of the relatively small dataset. The performance of the three best classifiers SVM-R, RF, and BG, aligns with the conclusions of [40] and [37]: they found these three binary classifiers to perform the best on 115 open-source datasets tackling a variety of real-world problems in medicine and biology (but not related to lameness detection). Although, on this dataset, the SVM classifier with radial kernel achieved the best performance in terms of accuracy and F1-score, it might not be the case for other datasets. This is a well-known machine-learning challenge, also known as the “no-free-lunch” theorem, that suggests that no algorithm can outperform all others for all problems [41]. Our recommendation would then be to try several classifiers, and the SVMs with radial kernel, random forests, and gradient boosting classifiers provide a good starting point.